### v2

# A Rigorous Behavior Assessment of CNNs Using a Data-Domain Sampling Regime

Abstract
We present a data-domain sampling regime for quantifying CNNs’ graphic perception behaviors. This regime lets us evaluate CNNs’ ratio estimation ability in bar charts from three perspectives: sensitivity to training-test distribution discrepancies, stability to limited samples, and relative expertise to human observers. After analyzing 16 million trials from 800 CNN models and 6,825 trials from 113 human participants, we arrived at a simple and actionable conclusion: CNNs can outperform humans and their biases simply depend on the training-test distance. We show evidence of this simple, elegant behavior of the machines when they interpret visualization images. osf.io/gfqc3 provides registration, the code for our sampling regime, and experimental results.
keywords:
Quantification, convolutional neural network, sampling, graphical perception, evaluation.
A toy example of a training data sampling regime. We quantify behaviors of convolutional neural networks (CNNs) in response to training inputs. Four data-domain sampling methods are used to supervise CNNs before sending them to take the same test. Introduction
A crucial consideration when using machine learning models, such as a convolutional neural network (CNN) or a large-language model, for visualization tasks is understanding when to trust their results and how to effectively train them [23]. Assessing their behavior quantitatively is essential for evaluating their reliability and identifying potential limitations. While most quantification studies emphasize algorithmic accuracy, highlighting vulnerability [18, 50], deficiencies in visual literacy [3, 48, 22, 17], and appearance-based robustness (e.g., changes in background colors or line widths [11]), evaluations explicitly addressing data-domain issues remain rare [53]. Haehn et al. [18] examined data-domain concerns, but their training, test, and validation datasets were restricted to the image ratios from Cleveland and McGill [8], which are essentially the test samples for human observers. To rigorously evaluate both unsupervised and supervised algorithms, the training data must be carefully defined, as its selection can give observers advantages or disadvantages when tested and compared [5].
Furthermore, the quantification of algorithmic performance can facilitate alignment with human observers’ ability. This is important as even for the most popular visualizations (e.g., bar charts and scatter plots), the ability to perform chart-reading tasks is considerably challenging to human observers who rely on the coordination of several mental processes, including rapid perceptual computations [8] with respect to a visualization schema [39], ability to perform abstraction [37], working memory [38], domain knowledge [52], and interpretation processes [46]. However, it is important to recognize that the behavior constraints of CNNs differ from those of humans [14].
The goal of this work is to fill in the knowledge gaps to critically quantify supervised discriminative CNNs by focusing on data-domain sampling. Especially, we study model behaviors by manipulating training-test data distribution shift to answer the following question: How do sampled training data values influence the model behaviors on the test set? Through multiple experiments, we assess algorithmic behaviors from several perspectives: models’ robustness to out-of-distribution data samples (sensitivity), resilience to limited input resources (stability), and when they surpass human (relative expertise).
Method. To systematically answer our research question and quantify CNNs’ behaviors in visualization contexts, we followed reliable tests used in human studies that focus on data targets such as bar heights [2, 45] or ratios [8, 20]. We then designed a training data sampling regime to enable direct behavior quantification during the test phase, after training models from shifting training-test data sampling distributions (see a toy example in A Rigorous Behavior Assessment of CNNs Using a Data-Domain Sampling Regime): (1) Independent and identical data distribution (IID): simple random sampling from the same data domain as the test set, as the baseline. (2) Coverage (COV): sampling values that are distant from each other in the training set; (3) Adversarial (ADV): sampling values that are positioned away from the test set values; (4) Out-of-distribution (OOD): sampling values concentrated in the left portion of the data domain. Through three controlled and two exploratory studies, we systematically quantify the reading behavior changes in response to these training data samples, after sending observers to take the “same” tests in each of the multiple runs, so that behaviors are comparable.
Results. We have new findings to interpret CNNs’ ability to decode charts. First, CNNs achieved higher accuracy and lower uncertainty with IID and COV in the ratio estimation task (section 4). Second, reducing the number of training samples increases both inference error and uncertainty, which explains the underperformed CNNs in Hanehn et al.’s experiment [18]. Third, an unexpected finding is the benefits of the coverage sampling COV: increasing training data span and evenly distributing samples within the span leads to higher stability than the most widely used IID-supervised models, especially when the total number of unique values is limited. Fourth, we also show an adversarial perspective when the training data are in the same domain as IID but introduced lower reliability. This means that CNNs, when used in visualizations, can easily be fooled (section 5). Finally, CNNs have much lower errors and uncertainty than people in some of the testing conditions (section 6). These results emphasize the importance of formalizing the training sampling domain of quantitative data when evaluating CNNs in visualizations.
Our work makes the following contributions: (1) A fundamental understanding of CNNs’ behaviors by focusing on quantitative data sampling unique to the visualization context; (2) A reusable controlled study method for comparing observers, either humans or CNNs. (3) New knowledge in CNNs’ behaviors through a set of experiments to quantify aspects of train-test discrepancies. We hope our research will both inspire new systematic AI behavior experiments in visualizations and support a cost-efficient interpretation strategy to understand complex algorithmic behaviors.
1 Background and Related Work
This section reviews the work that inspired ours: chart perception experiments in visualizations and quantification studies of CNNs.
1.1 Sampling Methods for Quantifying Visual Perception
Sampling is crucial in evaluating both human and algorithmic observers particularly in visualization tasks such as chart reading. In human studies, controlled sampling helps isolate the factors that influence reading. Cleveland and McGill in their 1984 seminal work (C-M) [8] sampled 7 out of 9 ratios ( row in Figure 1) derived from 10 heights: 10, 12.1, 14.7, 17.8, 21.5, 26.1, 31.6, 38.3, 46.4, 56.2 to spaced ratio samples seen by human observers. These samples were spaced evenly on a logarithmic scale to represent the power-law distribution of human length perception [24]. This work has been replicated and extended to crowdsourcing studies by Heer and Bostock [20], who chose a slightly different set of ratios ( row in Figure 1) from C-M’s original selection. Talbot et al. [45] adopted C-M’s ratios but instead sampled bar heights ( row in Figure 1) corresponding to those ratios.
Progress toward machine systems that quantify human-like understanding of visualizations has incorporated training data distribution as these human studies. For example, Haehn et al.[18] rounded Cleveland and McGill’s original 10 heights to the nearest integers, yielding 43 unique ratios (last row in Figure 1), which were randomly allocated for model training, validation, and test. This method assessed model inference accuracy under a specific human test data preparation paradigm, which did not consider that humans and models may have distinct decision mechanisms [14]. In this work, we explicitly separated training and test data to systematically assess algorithmic behaviors. This allows us to draw new conclusions and offer deeper insight into when and why CNNs succeed.
Type 1.
Adjacent bars
Type 2.
Aligned stacked bars
Type 3.
Separated bars
Type 4.
Unaligned stacked bars
Type 5.
Divided bars
1.2 Understanding Machine Learning Algorithm Behaviors
Our work is related to understanding the behavior of CNNs and aligns with three broad research directions. The first method concerns explainable approaches that introduce additional computational models [27] or visualizations [47, 55, 54, 43, 32] to interpret black-box behaviors. We avoided introducing new models, as this process itself is error-prone [41]. We also avoided visualization-based explainable methods (e.g., [50] and also see the excellent review [33, 6, 49]), as they are not quantitative [4] and may not be consistently perceived by human observers [15]. The second method involves human-algorithm alignment, where algorithms are used to explain human decision-making. This approach has only recently led to significant advances, offering a broad, global understanding of observers’ decisions [10, 35], primarily for categorical data in computer vision. The third method aims to trace algorithmic responses back to their training data by analyzing CNNs’ characteristics to interpret model behaviors [13, 51, 4]. This method has also been used in visualizations by posing different questions to an algorithm and observing its responses to interpret its behavior [3, 48, 22, 17].
In our work, we adopt the third approach to examine CNNs’ behaviors by probing their outputs through controlled alterations in input distributions. By analyzing the outputs in response to different training inputs, we aim to identify optimal strategies for training CNN models in a visualization context. To our knowledge, our method is the first attribution experiment designed to quantify CNN observers’ behaviors.
1.3 The Cost of Data Diversity
Producing large training sets is often a slow and data wrangling acquisition, and cleaning can be prohibitively expensive [25]. Therefore, we examine the reliability of the model under resource constraints to evaluate how models perform with varying numbers of unique inputs. An adversarial condition in vision is defined as “carefully crafted, seemingly normal inputs that have been slightly modified with imperceptible perturbations, causing the network to misclassify the image” [26, 16]. We define adversarial cases in the quantitative visualization domain as the training samples that share the same domain as test samples but may impair CNNs’ graphical perception ability. We leveraged an adversarial condition where the training samples are further from the test samples, and can limit the reliability of the models.
2 A Sampling Regime: Measuring Model Response to Training-Test Sampling Distribution Variations
This section presents the first contribution of this work: a training data sampling regime that quantifies model behavior under varying training-test discrepancies in the ratio or height data sampling domain.
2.1 The Problem Domain: Tasks and Visualization
We used a ratio estimation task, “judge what percent the smaller is of the larger” [8], because it required CNNs to reason about spatial relationships and had the highest inference errors among all tasks [18].
Ratio estimation task uses five bar chart types shown in Figure 2. Types 1–3 are bottom-aligned and thus judgment is by perceiving position along the common scale; and Types 4 and 5 must be compared by perceiving different bar lengths. Two target bars in each image are marked by a black dot and are adjacent to each other in all but Type 3. The ratio to be estimated is calculated as . In the human studies of position-length [8, 20, 45], people preferred position over length, and Types 1–3 outperformed Types 4–5.
2.2 The Target Data Sampling Domains
In our experiment, we used either the ratio or the taller bar height as the sampling target. Table 1 shows the range and count in the data domain.
Bar heights are integers: the taller bar height ranges from 6 to 85 (80 values), and the shorter bar height ranges from 5 to 84 (80 values), resulting in 3,240 unique () pairs. From these pairs, 2,081 unique ratios were derived, spanning from 5/85=0.059 to 84/85=0.988. We grouped these ratios into 94 evenly spaced bins (): , , …, and . For simplicity, each ratio bin is referred to by its midpoint value, such that , , …, , with . For Type 5, where two target bars are stacked on the same side, the constraint reduced the sampling domain slightly, limiting the ratio bins to . Figure 3 and Figure 4 illustrate the ratio samples and height samples along the axis, respectively.
| Ratio bin range | Types 1-4: | [0.06, 0.99] | (94) |
| (number of ratio bins) | Type 5: | [0.06, 0.98] | (93) |
| Taller bar height range | Types 1-5 | [6, 85] | (80) |
| (number of height samples) |
2.3 Problem Framing and Sampling Methods
Framing ratio estimation task for CNNs.
We frame the ratio estimation task as training a predictive model that can take an input image and output a predicted ratio .
Design principles.
We carefully controlled evaluation conditions, assessing a CNN’s ability and behavior by its responses to different sampling inputs by clearly defining the data domain. Given that models (observers) were compared in our experiments, we ensured a fair comparison by configuring conditions so that all observers took the same tests. In other words, test and validation sets were randomly chosen and held constant across different training conditions to ensure that models were evaluated using the same dataset for comparability. This control is crucial, as Cleveland and McGill [8] observed in their position-length experiment that human performance differed when judging ratios in different ratio bins. Because our focus is on data sampling, we keep appearance-based variables, such as line width, color, and gaps between bars, constant (See Apdx. Figure 26 for values used for these variables). The detailed steps below use ratio sampling as an example.
Step 1: Test and Validation Samples.
We selected = 19 ratio bins for each of the test and validation sets.
Step 2: Selection of Training Data Samples.
We assembled four sampling methods, each of which subsampled ratios from 56 ratio bins, i. e., half of the remaining 60% ratio bins. Here, could be set to any other numerical proportion of . We chose it for simplicity.
IID: The baseline sampling method used to train IID. Training and test samples follow the same simple random sampling.
COV: The distant-to-train condition used to prepare data to train COV. Samples were chosen sequentially by selecting that was the most distant from those that had been chosen so far. Thus, the first two selected were and , the minimum and maximum of available ratio bins. The third was the most distant ratio bin from and , and was close to the center of the sampling domain. This process continued by selecting the next as the most distant ratio bin from all previously selected ratio bins, until ratio bins were selected. In a sense, this sampling method encouraged the model to utilize the full data domain whenever possible.
ADV: The distant-to-test condition used to prepare training data to supervise ADV. The training samples were as distant from the test samples as possible. This is aimed to evaluate an adversarial condition, when the same number of samples as IID and COV were chosen from the same domain, thus looked similar, but distanced themselves from the test set. We can see that these samples form local clusters, as adjacent points had similar distances to the test set.
OOD: The out-of-distribution condition used to train OOD. We sorted all available 56 ratio bins in ascending order, and used the smallest 28 ratio bins to assemble a case of a biased training set. We called this OOD because the training set shifted to a different distribution from the test.
Step 3: Behaviors Under Limited Resources.
Given this sampling configuration, we evaluated stability under limited resources by reducing the number of training samples by half each time. Starting with of available ratio bins for training, we reduced this proportion to , , and , resulting in 14, 7, or 3 ratio bins, following the labeled order along the axis in Figure 3.
Rendering Images.
To generate an image, we first randomly select an in the training/validation/test set and then randomly select a () pair that yields a ratio within the selected bin, where and correspond to the shorter and taller target bar heights, respectively. For example, if was selected, both (19, 45) and (16, 38) are valid options for rendering the image. See Apdx. Figure 22 for the full set of () pairs for each ratio bin. The heights of other non-target bars in the image are assigned randomly. See Apdx. Figure 27 for examples of rendered images. We generated 60k/20k/20k images for training, validation, and test sets, respectively.
Model Configuration.
We used VGG19 [42] and ResNet50 [19], conducting five runs for each sampling configuration to simulate the process as if five individuals were making multiple independent decisions. The images and their corresponding ground truth ratios served as inputs to supervise the CNNs. Since the test set remained consistent across sampling methods, the model inferences were directly comparable. (See Appendix C for detailed model hyperparameter configurations.)
3 Summary of Studies
The following three sections detail our experimental methodology and findings. The studies are registered at osf.io/gfqc3.
Study I: Sampling effect (section 4): This study answers the question: How sensitive are CNNs to out-of-distribution samples? We controlled the sampling method, bar chart type, and CNN model architecture, evaluating performance using the mean absolute error. Our findings indicate that CNNs could achieve considerable accuracy, but were not robust to the distribution shift. Additionally, CNNs’ preference between position and length was minimal.
Study II. Stability (section 5): This study answers the question: How resilient are CNNs to limited input resources? We controlled the sampling method and downsampling level, assessing models based on mean absolute error and models’ intra-consistency. We found that downsampling the training set degraded model inference outcomes, and COV showed unexpected advantages that led to higher accuracy and reliability than IID.
Study III. Human-AI comparisons (section 6): This study answers the question: Can CNNs outperform humans? We controlled the observer (humans vs. CNNs) and bar chart type. Model performance was assessed using the midmean log absolute error for comparisons with Cleveland and McGill [8] and the mean absolute error otherwise. We also measure the human-CNN inter-consistency. Results show that CNNs had much lower errors and uncertainty than humans, and the errors of CNNs were also different from humans’.
4 Study I: Model Robustness to Distribution Shift
This section reports the second contribution of this paper: answering how models respond when training and test sets belong to different distributions. We conducted two experiments, each involving either ratio or bar height sampling, by varying three independent variables one at a time: the sampling method (4 levels), the visualization type (5 levels), and CNNs (2 levels) in a full-factorial design. The configuration led to runs in each experiment.
4.1 Design
4.1.1 Training Data Sampling
Exp 1 (ratio sampling robustness)
sampled the ratio bins as described in subsection 2.3, Figure 3.
Exp 2 (height sampling robustness)
applied the same sampling process to the 80 unique heights (Figure 4) as was used for ratio sampling, where 24/16/16 unique heights were in the training/validation/test sets applying the same ratio sampling process to the 80 bar heights.
4.1.2 Model Robustness Hypotheses
We had the following working hypotheses:
Sampling effect.H1 (Accuracy and Adversarial hypotheses). Models will achieve high accuracy with IID and COV sampling, but not with ADV and OOD sampling.
Sampling effect.H2 (Out-of-distribution hypothesis). Models trained in OOD will have reduced accuracy in regions where the training and test sets do not overlap.
Sampling effect.H3 (Chart type hypothesis): We will not observe the effect of visualization types on inference accuracy.
Several reasons led to these hypotheses. Both IID and COV encouraged diversity, with IID being commonly used due to its ability to produce accurate inferences. COV also increased diversity by enforcing a greater separation between training samples. We did not expect ADV and OOD to perform well on average, as ADV represented an adversarial case, and OOD involved a distribution shift in the sampling domain. For the same reason, in H2, we did not expect accuracy gains for test values that were absent from the training data domain. H3 could be supported because CNNs were largely affected by data sampling input, rather than human perception limits.
4.2 Analysis Method
Each CNN model produced 20K inferences. Given the large number of inferences, we made several decisions when analyzing the data. First, we aggregated predictions by () pairs, treating each aggregation as an independent trial for downstream statistical analysis. This was necessary to avoid false discoveries resulting from the large number of predictions. Second, the error metric was mean absolute error (MAE), defined as , consistent with the evaluation metric used in human observer experiments by Talbot et al. [45]. We did not use the logarithmic mean as in the C-M study [8] in this and the next sections, because taking the logarithm could disproportionately emphasize smaller errors. We also did not use mid-mean to trim data in these first two studies for practical considerations: we were interested in CNNs’ behaviors in real-world uses, where determining when to trim data may not always be feasible.
We examined the main effect of independent variables on prediction errors and performed statistical analysis using the General Linear Model (GLM). When a significant main effect was observed, we performed a post-hoc analysis using Tukey’s Honest Significant Difference (HSD) test to assess the significance of differences across the levels in a treatment condition. Finally, we calculated the effect size using Cohen’s [9] to measure the practical significance: “no effect” (), “small effect” (), “medium effect” (), and “large effect” ().
4.3 Results
| Variable | HSD | |||
| Study I Exp 1: Ratio sampling | ||||
| Sampling (VGG19) | ||||
| Type (IID, VGG19) | ||||
| Model (IID) | ||||
| Study I Exp 2: Height sampling | ||||
| Sampling (VGG19) | ||||
| Type (IID, VGG19) | ||||
| Model (IID) | ||||
Overview. We collected a total of 4 million predictions from 200 models in each experiment (). Figure 5 and Table 2 show the mean errors and the summary statistics. See additional analyses in Apdx. Figure 28 and Apdx. Table 6. Sampling method is a significant main effect. Sampling effect.H1 was supported for height sampling but not for the ratio sampling experiment. For the ratio sampling experiment, ADV (0.013) was in the same Tukey group as COV (0.011) and IID (0.011) (Figure 5 ), meaning that we did not observe a statistical difference among them. One reason for the lack of significance for ADV could be that it contained sufficient unique samples across the data domain.
For height sampling, COV (0.015) and IID (0.027) had lower errors than ADV (0.042) and OOD (0.059) (Figure 5 ). Now ADV was in a different Tukey group (Table 2). We suspected that the differences between IID, COV, and ADV conditions were caused by errors introduced by the smaller bar heights in Figure 6.
Striking generalization failure on OOD. Sampling effect.H2 was supported. To study the CNNs’ behaviors, we analyzed case-by-case CNNs’ inference accuracy by ratio bins (Figure 6 ) and by heights (Figure 6 ) computed from the two experiments correspondingly. In Figure 6 , the diagonal line means that the prediction is a perfect match to the ground truth, and the closer to this line, the smaller the inference errors. Figure 6 shows the mean absolute errors against bar heights, lower error indicates better model performance. We observe that knee-shaped curves emerge in the out-of-train regions in OOD in both ratio sampling and height sampling experiments, showing errors that increase almost linearly. Interestingly, the in-train test data points in OOD maintain low errors (0.008 for ratio sampling and 0.025 for height sampling). This result may support that CNNs’ inference accuracy depends on the coverage of input: IID did not guarantee sampling of the smallest or the largest values compared to COV and the minimum values of OOD. Our experiments in the next section validate this conjecture by stress-testing CNNs with limited training samples.
CNNs had limited ability to read smaller bar heights. We
performed Tukey’s HSD test on the 80 bar height levels and found that smaller bar heights (H15 pixels) increased inference errors compared to larger bar heights, independent of sampling methods. This small height effect was also observed in the ratio sampling experiment, where having H pixels was necessary to lower the inference errors (See Apdx. Figure 29).
The effect of types varied by experiments. Sampling effect H3 is partially supported. Type was a significant main effect in ratio sampling only. We see that CNNs’ differences in inference accuracies between types are much smaller than those of humans.
Model was also a significant main effect. Specifically, VGG19 outperformed ResNet50 in our experiments. We suspect that this performance difference can be attributed primarily to the way we configure the network, rather than to any inherent advantages or disadvantages of these architectures. Given these findings, we focus our subsequent discussion on results obtained for Type 1 and VGG19, and leave other types and ResNet50 in Apdx. Figure 28.
5 Study II: Downsampling to Test Model Stability
This section reports results of the sampling stability, which examines how the number of unique training samples impacts inference errors.
Understanding model stability can help reduce training costs and guide the selection of more resilient sampling strategies. We also ran two experiments, each downsampling ratio bins or bar heights. We varied two independent variables, one at a time: the sampling method (4-levels) and the downsampling level (3-levels) in a full-factorial design. The configuration led to conditions for each experiment, and running 5 times led to a total of models in each experiment. Besides accuracy, we quantified intra-consistency to measure the variability of the inference results obtained using the same sampling method at different downsampling levels. Higher consistency indicates better model stability.
5.1 Design
5.1.1 Training Data Sampling
We used the same test sets as Study I. The implementation of downsampling was straightforward, described in subsection 2.3. The training samples were a subset of those used in Study I, section 4. The number of training ratio bins was 14, 7, or 3, and the number of heights was 12, 6, or 3, at downsampling levels of 15%, 7.5%, or 3.75% respectively, taken in order from our sampling configurations in Figure 3 and Figure 4.
5.1.2 Model Stability Hypotheses
We had the following working hypotheses:
Stability.H1 (downsampling accuracy hypothesis): Decreasing the number of unique samples in the training set will increase CNNs’ inference errors.
Stability.H2 (downsampling stability hypothesis): The four sampling methods will not be equally robust.
These hypotheses could be supported because COV and IID remained diverse, while OOD could introduce greater distribution shifts towards the smaller values. Under constrained resources, ADV also introduced out-of-distribution cases, further reducing its reliability.
| Variable | HSD | |||
| Study II Exp 1: Ratio downsampling | ||||
| Level | ||||
| 15% | ||||
| 7.5% | ||||
| 3.75% | ||||
| Study II Exp 2: Height downsampling | ||||
| Level | ||||
| 15% | ||||
| 7.5% | ||||
| 3.75% | ||||
5.2 Analysis Method
We used the same method to analyze inference errors as described in Study I. In additionally to mean absolute error, we further measured the intra-consistency of model inferences between downsampling levels by the Pearson correlation coefficient [9]. A higher Pearson correlation indicates greater intra-consistency and stability for a sampling method in response to limited resources.
5.3 Results
Overview. We collected a total of 1.2 million predictions () for each of the ratio and height stability experiments. Figure 8 and Table 3 show the mean errors and the summary statistics. Both hypotheses were supported. (More analyses in Apdx. Figure 30.)
Reducing the number of unique training samples increased inference errors, with COV exhibiting half the errors of IID. Stability.H1 was supported. The downsampling level was a significant main effect on errors. In the ratio sampling experiment, inference errors for IID increased more than tenfold, from 0.011 at a downsampling level of 30% to 0.146 at 3.75%. For COV, errors rose about sevenfold, from 0.010 to 0.073. COV and IID were in the same Tukey group until the downsampling condition (Table 3). At downsampling level, ADV (0.051) was in the same group as IID (0.011) and COV (0.010), but its error increased sharply () between 7.5% and 3.75%. These results suggested that COV was the most stable sampling method among the four. The same trend was observed in the bar height experiment, where COV again produced the lowest errors (Figure 8).
COV is the most stable sampling method. Stability.H2 was supported. We assessed model stability using intra-consistency, calculated as the correlation of model inferences across different downsampling levels. As shown in Figure 9, IID and COV had the largest intra-consistency () between the 30%, 15%, and 7.5% levels in the ratio downsampling experiment. For IID, consistency dropped to when including the 3.75% level, whereas COV maintained a high consistency of . Intra-consistencies were much lower for ADV () and OOD () on average. In the height downsampling experiment, COV also showed the highest intra-consistency (), compared to IID (). ADV () had even lower consistency than OOD (). These results indicate that the COV sampling method is the most stable across different downsampling levels, maintaining high intra-consistency even at extreme reductions.
Figure 7 provides a case-by-case analysis of the ratio downsampling experiment, further illustrating the stability of COV. OOD errors increased almost linearly as test cases moved further out-of-distribution. Both IID and ADV also exhibited OOD-like behavior when training samples were limited. In contrast, COV enforced in-distribution conditions for all test cases, preventing errors from rising linearly; instead, they remained bounded by the evenly distributed training samples. This is most evident at level 3.75%, where only three samples were used for training. The errors of COV were stable, and lower errors occurred when test samples were close to the training samples—specifically, at the leftmost, rightmost, and middle positions.
6 Study III: Human vs. CNN: A Comparative Study
Given the relatively small errors in the first two studies when applying IID and COV, we proceeded to compare CNNs and humans and answer the question: to what extent do CNNs and humans differ? For the human observer, the bar chart type was the only independent variable considered. For the CNN observer, we introduced a new baseline condition (IID-large) that utilized the entire remaining 60% of ratio bins or heights without downsampling. This corresponded to 56 ratio bins in the ratio sampling experiment or 48 height bins in the height sampling experiment. We denoted this condition as “IID-large”. We also reused the IID and OOD conditions from Study I section 4. We did not use COV since its error was bounded by the error of IID.
6.1 Design
6.1.1 Human Experiment Data Collection
We adapted the position-length experiment in Cleveland and McGill [8] using the 10 integer bar heights of . Exhausting all pairs generated 45 height combinations, resulting in 37 unique ratios across 23 ratio bins: 0.18, 0.26, 0.27, 0.38, 0.39, 0.45, 0.46, 0.47, 0.48, 0.55, 0.56, 0.57, 0.58, 0.67, 0.68, 0.69, 0.7, 0.8, 0.81, 0.82, 0.83, 0.84, 0.85.
We conducted the experiment on the LabIntheWild [40] crowdsourcing platform (see the experiment user interface in Apdx. Figure 31). Participants were instructed to make a quick division between the heights of the shorter marker bar and the taller marked bar. They first completed 5 training trials—one for each bar chart type—with immediate feedback provided after each response. After that, each participant completed 25 trials, consisting of 5 chart types with 5 trials per type in random order.
6.1.2 Human-Machine Comparison Hypothesis
We had the following working hypotheses:
Human-AI comparisons.H1 (accuracy hypothesis): Our CNNs’ errors will be significantly lower than humans’.
This hypothesis could be supported based on our observation comparing our results in the first two experiments to those in the human experiments of Cleveland and McGill [8]. CNNs appeared to have much smaller errors compared to those of human observers.
6.2 Analysis Method
We used mid-means of the log absolute error (MLAE) [8] to measure accuracy, in order to compare our results to those published in literature, i. e., Cleveland and McGill [8], Heer and Bostock [20], and Haehn et al. [18]. Here . The height sampling experiments used mean absolute error to compare to that of Talbot et al. [45]. Finally, we evaluated human-CNN inter-consistency using Pearson correlation analyses.
6.3 Results
| Variable | HSD | |||
| Study IV Exp 1: Ratio sampling | ||||
| Observer | ||||
| Type (human) | ||||
| Study IV Exp 2: Height sampling | ||||
| Observer | ||||
| Type (human) | ||||
Overview. We collected human observers’ results from 1,575 trials from 63 participants, who spent an average of 10.0 seconds per trial. Figure 10, Figure 11, and Table 4 show the mean errors and the summary statistics. Overall, our human experiment results showed slightly higher error than Cleveland and McGill’s (Figure 10 ), but the trend across the five chart types remained consistent with the findings of Cleveland and McGill [8] and Heer and Bostock [20] (Figure 10 ). (Apdx. Figure 32 has more details.) CNNs had lower errors than humans. Human-AI comparisons.H1 was supported. The observer was a main effect. For ratio sampling, human errors (0.056) were about four times that of CNNs’ IID-large (0.013) (Figure 10 ), and for height sampling, human errors (0.058) were about three times of CNNs’ IID-large (0.019) (Figure 11 ).
Chart type was also a significant main effect. The differences between types were observed in human experiments (Figure 10 and Figure 11 ). In addition, CNNs had smaller uncertainty (95% CI for Type 1, Baseline, ratio: 0.002, height: 0.002) compared to human observers (95% CI, ratio: 0.007, height: 0.004).
Human-CNN inter-consistency is lower than the intra-consistency between CNNs (Figure 13). In ratio sampling: ; and in height sampling: . This suggested that human errors differed from those in CNNs.
7 General Discussion
This section presents explanations for our findings, reflects on approaches for quantifying CNNs’ capabilities, and provides practical recommendations for training data preparation.
7.1 The Effect of Training-Test Distances and Reusability
One of the most intriguing findings in our study was the stability of the COV sampling method, particularly when the total number of training samples was limited (Figure 8). We hypothesize that this stability stems from the key property of COV: it ensures a broad coverage of the data space while maintaining an even distribution of training samples. By maintaining a well-spread distribution of training samples, COV effectively minimizes large gaps in the training space, ensuring that test samples are likely to have nearby training samples for reference.
To validate this hypothesis, we analyzed the relationship between inference error and the training-test distance—the average distance from each test sample to its closest training sample. As shown in Figure 14, there was a strong positive correlation in both ratio () and height () experiments, confirming that models performed more accurately when test samples were closer to their nearest training neighbors. Among all sampling strategies evaluated, COV consistently exhibited the lowest training-test distance at equivalent downsampling levels, further supporting its advantage in stability.
To examine the reusability of this stability result, we investigated whether the impact of training-test distance extends beyond the ratio estimation task. We observed a similar pattern in two additional challenging real-world counting tasks. The first is counting cells in fluorescence-light microscopy images rendered using the method of Lehmussola et al. [29]. It required CNNs to identify cells based on continuous color, textures, and shapes. The second is counting network nodes in node-link diagrams rendered using high-fidelity graph simulations by Lancichinetti et al. [28]. We manipulated the community size and required CNNs to perceive the shape of the diagram to make a prediction. The results in Figure 15 suggest that the effect of training-test distance can be generalized to new tasks. (Apdx. subsection D.3 has experiment configuration details.)
7.2 Implications to Training Data Preparation Concerning Reliability, Stability, and Cost-effectiveness
The results of our study suggest that the training sampling choices were a significant main effect on CNNs’ inference errors. Using a few human ratio samples may disadvantage CNNs, as exposure to a wider range of ratio samples could improve their capabilities. Implicitly mixed sampling conditions in the training could potentially increase inference uncertainty, because inference errors from an OOD-trained model could be much larger than those from an IID-trained model. On the other hand, conditions that benefit CNNs could also appear. When training and test samples are close neighbors, the learning algorithm can gain an advantage, leading to higher accuracy.
From a reliability perspective, we found that CNNs could achieve considerable accuracy. However, they are not robust to the train-test distribution shift. In-distribution tests remained accurate, but errors increased linearly for out-of-distribution cases (Figure 6). From a stability perspective, models should try to cover the data range as much as possible, especially when resources are limited or when sampling more data is too expensive (Figure 14). Under limited resources, the most commonly used IID sampling method does not guarantee stability (Figure 9). The lack of robustness in the ADV condition also suggests that chart reading CNNs could also be easily fooled, which could lead to serious consequences for high-stakes domains. From a cost-effective perspective, we found that using COV can substantially reduce the training time: on average 69.7%, 82.6%, and 96.4% less time compared to IID, ADV, and OOD, respectively (Apdx. Figure 37).
These results imply that we should always include more samples and ensure a broader training coverage when possible. This practice helps prevent adversarial conditions, in which training and test samples share the same domain but are distant from each other, a scenario that can severely degrade the accuracy of CNNs. Finally, we suggest enhancing transparency to mitigate misuse and misinterpretation. A practical measure is to incorporate a model card [36], providing information on how training samples are selected, along with their ranges and distributions, image size, and the number of unique samples.
7.3 Comparison to the Results of Haehn et al. [18]
Haehn et al. [18] stated, “In other experiments, such as the position-length experiment, CNNs cannot complete the task.” However, our Study III in section 6 demonstrated that CNNs can outperform humans. We provide an explanation for this mismatch from a data sampling perspective. In our experimental setup, the heights of target bar samples ranged in [5, 85] pixels, whereas Haehn et al. followed Cleveland and McGill’s approach, using only 10 discrete heights. As a result, our configuration led to a substantially larger set of unique () pairs (see Apdx. Figure 22). Although the smaller training-test distance in Haehn et al.’s experiment (as shown in Figure 1) provided great advantages to CNN observers, their limited variation in training () pairs was among the reasons that hindered models’ inference accuracy.
From the attribution study perspectives, their approach of random assignment of the C-M test samples into the training, validation, and test sets made it hard to analyze the influence of the training set on large errors and uncertainty in model inference outcomes. Their design and analyses may have introduced large variations in test conditions as well (see Apdx. Figure 23). In contrast, our study explicitly controlled training and test sets, providing an interpretable perspective.
These sampling methodological differences highlight the impact of sampling strategies on algorithmic performance. Our results aligned with computer vision studies that CNNs’ behavior constraints differ from humans’ [14]. These spatial tasks seem to be manageable for CNNs when we have a good knowledge of data preparation. Directly adopting experimental designs from human studies may introduce a methodological bias, leading to models that may not generalize well [12]. Therefore, comparing models trained under such methodological biases against humans, especially without a fixed and identical test set, may not constitute a fair evaluation.
7.4 The effect of Small Stimuli
Our findings from Studies I and II showed that small bar heights ( pixels for height sampling and pixels for ratio sampling) led to increased errors for IID and ADV, but not for OOD (Figure 6). The first explanation was that OOD encountered more images with short bars during training, as its smaller sampling domain resulted in a smaller training-test distance within the sampling regions. A second explanation could be the effect of sideness: we only selected the left side with small bar heights in our experiments and did not test OOD using regions with the largest bar heights. If similar errors occurred in the larger-side sampling, it would suggest that the error was not due to bar height itself but rather a consequence of our design choice related to the sideness. We thus examined the other side of the bar height or ratio samples under OOD conditions to train models. We observed an asymmetry in the height sampling conditions (Figure 16), suggesting that smaller heights indeed led to larger errors.
This behavior contrasts with human perception that follows Weber’s Law [24], making ratio judgments independent of the absolute magnitude of the stimuli. Apdx. subsection D.1 provides explanations to the short height effect from CNNs’ architectural perspective. To improve this condition, one can increase the input image size so that smaller bar heights are proportionally increased, or use networks and modules that are designed for small-size elements [30, 31].
7.5 Limitations and Future Work
In this work, we designed new attribution study method to quantify algorithmic behaviors, by explicitly considering training-test discrepancies in four sampling strategies. There are other sophisticated sampling methods, e.g., those following target data distribution [1] could also be more effective that are worth future investigation. We’ve developed a new attribution study method that quantifies algorithmic behaviors, offering foundational scientific insights into effective training data sampling for visualization tasks. While this method advances our understanding of algorithmic quantification, it remains a stepping stone, not a standalone solution for real-world challenges. Addressing the intricate contexts and patterns of practical analytical tasks will necessitate a series of future studies to effectively leverage our quantification methods in quantifying observers. Finally, fundamental theories to quantifying CNNs remain scarce. Methods to test the generalizability of the findings to other visualization types, beyond those tested here, are needed to study the alignment between humans and machines specifically designed for visualization uses.
8 Conclusions
Our sampling regime reveals fundamental insights into CNNs’ graphical perception: while models achieve superhuman accuracy in controlled settings, their brittleness under distribution shifts, especially the adversarial sampling (ADV), suggests the need for rigorous training-test separation to test CNNs in visualization uses. Coverage-based sampling (COV) emerged as a robust strategy, largely due to its maximized data span and evenly distributed samples, which is crucial for reducing training-test distance without prior knowledge. Smaller visual stimuli (e.g., short bars) were particularly error-prone, suggesting that training protocols must account for mark size. Using our sampling regime, we have disentangled the effects of sampling strategies, emphasizing the need for careful experimentation to understand CNN model behaviors. This understanding is critical when human observers rely on automated tools in scenarios involving human-AI collaboration—we need to know when to trust AIs and when not to.
Our research findings have yielded new insights into these black-box CNNs’ abilities to interpret charts.
Quantification Method: it is more informative to report case-by-case accuracy and beyond-average performance metrics. Notably, even in OOD conditions, in-sample data still performed well.
Efficiency: Prioritize diverse sampling over large datasets—COV requires fewer samples than IID for comparable accuracy.
Transparency: When deploying CNNs in real-world visualization tasks, model cards should include details on sampling methods, such as data span and train-test distributions.
Visual encoding: Sufficiently large visual marks may be necessary to reduce errors.
Beyond the recommendations we have made, we hope our sampling regime and the probing methods provide the community with a means to quantify behavior differences between humans and machines.
Acknowledgements.
We thank Justin Talbot, Vidya Setlur, and Anushka Anand for sharing their results, which we have used for human-CNN comparisons in section 6. We also thank all participants who volunteered in our online experiment for their time and effort. Finally, we thank the reviewers for their feedback, which helped us improve the quality of this paper.Image Copyrights
We, as authors, state that the plots in this paper are and remain under our own copyright, with the permission to be used here. We have also made them available at osf.io/gfqc3 and Google Drive.
A Rigorous Behavior Assessment of CNNs Using a Data-Domain Sampling Regime
Additional material
Appendix A Reproducibility & Access to Code/Models/Data
In this Appendix, we report experimental details for human and CNN experiments, reported in this paper. Experiment results and code can be accessed at osf.io/gfqc3 or Google Drive. The human behavioral dataset contains 1,575 trials across 63 observers for the ratio sampling experiment, and 4,903 trials across 50 participants for the height sampling experiment.
Appendix B A Comparison of Sampling Input Distributions and Data Processing Approaches
The sampling values used in the corresponding experiments are marked in Figure 22 and sampling values are chosen below.
Cleveland and McGill [8] ( in Figure 22). Seven ratios from a geometric sequence evenly spaced on a log scale of 10 bar heights , where (). The 45 combinations of those 10 heights led to 9 distinct ratios (0.18, 0.22, 0.26, 0.32, 0.38, 0.46, 0.56, 0.68, 0.83). 0.22 and 0.32 were not used because they were too close to 0.18 and 0.26.
Heer and Bostock. [20] used the following ratios: 0.18, 0.22, 0.26, 0.32, 0.47, 0.55, 0.56, 0.69, and 0.85, with unknown bar height choices.
Talbot et al. [45] ( in Figure 22) drew bar charts on a 200px200px canvas. Their three taller bar heights were 50px, 100px, and 150px. Ratios were the same 7 ratios as Cleveland and McGill’s [8]: 0.178, 0.215 0.261, 0.316, 0.383, 0.464, 0.562, 0.682, 0.825.
Haehn et al. [18] ( in Figure 22). Used the 10 integer bar heights (10, 12, 15, 18, 21, 26, 32, 38, 46, 56). The simulation of Haehn et al.’s 12 exemplar runs in Figure 23.
Our replication also used the 10 integer heights (10, 12, 15, 18, 21, 26, 32, 38, 46, 56). We also removed height combinations [(10, 46), (12, 56)] and [(10, 32), (12, 38), (15, 46), (18, 56)], which will produce ratios 0.18 and 0.32 in the original Cleveland and McGill’s experiment if the heights were not rounded. in Figure 22 except for the excluded combinations near ratio 0.18 and 0.32 were used in our experiment.
Appendix C CNN Model Configuration
We trained both VGG19 and ResNet50 on multiple NVIDIA P100 GPUs, using CUDA 11.2, implemented using TensorFlow 2.4 and Python 3.9. We reused the model configuration in Haehn et al.’s [18]. We briefly describe the configuration here to make this paper self-containable. We used VGG19 or ResNet50 as the feature extractor, followed by a one-hidden-layer multilayer perceptron (MLP) for regression. The hidden layer had 256 neurons, with ReLU as activation function and 0.5 dropout to prevent overfitting. We trained CNNs from scratch for at most 100 epochs using mini-batch stochastic gradient descent (SGD) with a batch size of 32. Training stopped early when validation loss did not improve for ten consecutive epochs. The learning rate was 0.0001 and we set Nesterov momentum [44] to 0.9 to accelerate convergence by leveraging past gradients and reducing oscillations, making updates smoother and more stable. During training, we used mean square error (MSE) as the loss function.
Appendix D Additional Study Details
D.1 Architectural Influences on Model Accuracy
In subsection 4.3 we found that VGG19 outperformed ResNet50 in terms of accuracy. This disparity may stem from VGG19’s architecture. With its regular stacking of convolutional layers, is well-suited for capturing the local patterns and structural features in the images, such as the bar heights and the marked dot. In contrast, ResNet50’s deeper architecture and residual connections, while powerful for semantic information, might introduce unnecessary complexity or dilute the focus on localized features in this simpler setting.
We also observed a small bar height effect in subsection 4.3, which may be explained by how CNNs are implemented. CNNs (e.g., VGG19) are optimized for classification tasks, which use pooling layers to reduce spatial resolution. While this helps with object recognition by creating abstract features, it causes crucial information from small features, like short bars, to be lost in deeper layers [34].
D.2 Additional Details about the Human Experiment
Ratio and Height Sampling
We used the same 10 heights as Cleveland and McGill, rounded to the nearest integer: 10, 12, 15, 18, 21, 26, 32, 38, 46, and 56. In Cleveland and McGill’s original experiment, the 45 combinations of these 10 heights yielded 9 ratios, of which two ratios (0.22 and 0.32) were not used. In our experiment, since the heights were rounded, we obtained 9 ratio groups, which was the same as Haehn et al. shown in Figure 1 in the main text. Unlike Haehn et al., we also excluded the ratio groups corresponding to the ratios 0.22 and 0.32 that Cleveland and McGill omitted, preserving 39 height pairs in 7 ratio groups.
Image Generation and Rendering
We drew bar charts on 100100 pixel resolution saved in a vector format to ensure proper display at various screen resolutions on participants’ devices. The minimal monitor resolution allowed for the experiment is 720P. The image height was scaled to either one-third of the browser window height or 300 pixels, whichever was larger.
Experiment Procedure and Instruction
Prior to beginning the experiment, participants filled in a demographic questionnaire, and their gender, age, education level, and nationality were collected. Participants were trained on the ratio estimation task. To familiarize participants with the task requirements, they completed 5 practice trials, one for each chart type, using randomly generated heights distinct from the 10 heights used in formal experiment. The practice trial interface is illustrated in Figure 31. During practice, participants were given immediate feedback after making their judgments. In the formal experiment, for each chart type, the 5 trials were randomly selected from the 7 possible ratio groups. Each participant completed 25 trials, consisting of 5 trials for each of the five bar chart types, in random order. The trials were presented with all chart types intermixed. No feedback was provided during the formal experiment. Panning and zooming were not allowed. Participants filled in their answers in the blank, and clicked the “Next” button to proceed to the next task.
Outlier Removal
Seven out of 63 participants were outliers and excluded from the study, as their errors exceeded three times the interquartile range.
D.3 Additional Configuration Details for Reuse
We measured whether the effect of training-test distance was generalizable (main text subsection 7.1). We introduced two visualization types: continuous color and texture and node-links for two new counting tasks: cell counting and node counting, where data were generated in high-fidelity simulations. We selected these two tasks not only because they represent classic problems in biology and network analyses, but also because they utilize new visualization types distinct from the position/length in ratio estimation task: texture atop of shapes and node-links. Akin to sample ratios for bar charts, here we sampled cell counts and node counts. The cell counts ranged from 80 to 159 (Figure 35), and the node count from 20 to 99 (Figure 36), both yielding 80 unique samples. We followed the same proportion configuration used in the ratio estimation task to configure the training, validation, and test sets: 20-20-30 (16-16-24 samples) for test-validation-training, after applying the four sampling methods to sample the sets. The downsampling experiment reduced the number of unique training samples to 15%, 7.5%, and 3.75%, or 12, 6, and 3 samples. In both cell and node counting experiments, we generate 6K/2K/2K images for training, validation, and test sets, respectively.
| VGG19 | ResNet50 | |||||||||
| Experiment (effect) | HSD | HSD | ||||||||
| Study I Exp 1: Ratio sampling | ||||||||||
| Type (all sampling) | ||||||||||
| Type (IID) | ||||||||||
| Type (COV) | ||||||||||
| Type (ADV) | ||||||||||
| Type (OOD) | ||||||||||
| Sampling (all types) | ||||||||||
| Sampling (Type 1) | ||||||||||
| Sampling (Type 2) | ||||||||||
| Sampling (Type 3) | ||||||||||
| Sampling (Type 4) | ||||||||||
| Sampling (Type 5) | ||||||||||
| Categorical (Type 1) | ||||||||||
| Categorical (Type 2) | ||||||||||
| Categorical (Type 3) | ||||||||||
| Categorical (Type 4) | ||||||||||
| Categorical (Type 5) | ||||||||||
| Study I Exp 2: Height sampling | ||||||||||
| Type (all sampling) | ||||||||||
| Type (IID) | ||||||||||
| Type (COV) | ||||||||||
| Type (ADV) | ||||||||||
| Type (OOD) | ||||||||||
| Sampling (all types) | ||||||||||
| Sampling (Type 1) | ||||||||||
| Sampling (Type 2) | ||||||||||
| Sampling (Type 3) | ||||||||||
| Sampling (Type 4) | ||||||||||
| Sampling (Type 5) | ||||||||||
| Categorical (Type 1) | ||||||||||
| Categorical (Type 2) | ||||||||||
| Categorical (Type 3) | ||||||||||
| Categorical (Type 4) | ||||||||||
| Categorical (Type 5) | ||||||||||